How many quarter are there altogether in 11 1/2
1
Expert's answer
2011-08-17T07:28:20-0400
To solve this question we must divide 11 1/2 by 1/4
11 1/2 = 23/2
(23/2) / (1/4) = 23/2 *4 = 46
Answ: there are 46 quaters in 11 1/2.